19.50 ml. of a mixture of NH3 and H2 was completely decomposed by sparking into nitrogen and hydrogen. 40 ml of oxygen was then added and the mixture was sparked again. After cooling to the room temperature the mixture was shaken with alkaline pyrogallol and a contraction of 6 ml. was observed. Calculate the % of NH3 in the original mixture. (Assuming that nitrogen does not react with oxygen.)

Grade:

1 Answers

Sunil Kumar FP
9 years ago
nh3 ---1/2 n2 + 3/2h2
let the volume of nh3 be a
volume of n2=a/2 and h2=3/2a
total volume of h2=3/2a +50 -a=50+a/2
h2+ 1/2o2-----h20
volume of o2 absorbed=6=40-(50+a/2).5
a=36
percentage of nh3=36/50*100=72%
thanks and regards
sunil kr
askIITian faculty